OpenStudy (anonymous):

Solve the inequality algebraically. (x+4)/(x-2) < 1

OpenStudy (anonymous):

\[(x+4) / (x-2) \le 1\]

OpenStudy (helder_edwin):

u have \[ \large \frac{x+4}{x-2}\leq1 \] then \[ \large \frac{x+4}{x-2}-1\leq0 \] do u get this?

OpenStudy (anonymous):

Yep, I do. I don't know what to do afterwards.

OpenStudy (helder_edwin):

now do the substraction of what is on the left tell me what u get

OpenStudy (anonymous):

6 / x-2

OpenStudy (helder_edwin):

great now 6>0 right?

OpenStudy (anonymous):

Yup.

OpenStudy (helder_edwin):

so for the fraction to be \(\leq0\) then x-2 has to be positive or negative?

OpenStudy (helder_edwin):

so?

OpenStudy (helder_edwin):